# A Review of Graph Neural Networks in Epidemic Modeling

### v2

@@ Abstract.
Since the onset of the COVID-19 pandemic, there has been a growing interest in studying epidemiological models. Traditional mechanistic models mathematically describe the transmission mechanisms of infectious diseases. However, they often fall short when confronted with the growing challenges of today. Consequently, Graph Neural Networks (GNNs) have emerged as a progressively popular tool in epidemic research. In this paper, we endeavor to furnish a comprehensive review of GNNs in epidemic tasks and highlight potential future directions. To accomplish this objective, we introduce hierarchical taxonomies for both epidemic tasks and methodologies, offering a trajectory of development within this domain. For epidemic tasks, we establish a taxonomy akin to those typically employed within the epidemic domain. For methodology, we categorize existing work into Neural Models and Hybrid Models. Following this, we perform an exhaustive and systematic examination of the methodologies, encompassing both the tasks and their technical details. Furthermore, we discuss the limitations of existing methods from diverse perspectives and systematically propose future research directions. This survey aims to bridge literature gaps and promote the progression of this promising field, with a list of relevant papers at https://github.com/Emory-Melody/awesome-epidemic-modeling-papers. We hope that it will facilitate synergies between the communities of GNNs and epidemiology, and contribute to their collective progress.
@@ 1. Introduction
Epidemiology has long been a critical field, with its origins tracing back to ancient societies that observed patterns of disease spread (bramanti2012ancient, 1, 2). Although the conceptualization of epidemiology has evolved over time, the terms health and control have been predominantly associated with it since 1978 (frerot2018epidemiology, 3). Currently, the World Health Organization (WHO) describes epidemiology as the investigation into the distribution and determinants of health-related states or events, encompassing a broad spectrum of issues including disease transmission, vaccination efforts (fine2015another, 4, 5), cancer and diabetes treatment, etc. This definition underscores the field’s emphasis on controlling health-related issues and making informed decisions. A pertinent illustration of this is the COVID-19 pandemic, which rapidly infected millions worldwide, placing immense strain on the production and distribution of medical resources (cm2020does, 6, 7). Decision-making processes like allocation of resources, are crucial in mitigating the impact of diseases and saving lives, highlighting the importance of advancements in epidemiology (emanuel2020fair, 8).
@@
To address a range of health-related challenges, there is an indispensable need for epidemic modeling, and researchers have devised various mechanistic models (funk2018realtime, 9, 10). These models, grounded in mathematical formulations, simulate the dissemination of infectious diseases by incorporating biological and behavioral underpinnings. By considering factors such as population, they yield insights into patterns of disease transmission and the efficacy of intervention strategies, thereby playing a pivotal role in shaping public health policies (louz2010emergence, 11, 12). However, these knowledge-driven methods often depend on oversimplified or fixed assumptions that can lead to biases in modeling. Consequently, this compromises both the accuracy of predictions and their ability to generalize across different contexts.
@@
To overcome the limitations of mechanistic models, there is an emerging trend to adopt data-driven approaches in epidemic forecasting tasks, with a particular emphasis on machine learning and deep learning models (shorten2021deep, 13). Specifically, Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) have demonstrated great success in various epidemiological predictive tasks, including forecasting daily new case counts, estimating virus reproduction and doubling times, and determining disease-related factors (wu2018deep, 14, 15, 16). Despite their effectiveness in these tasks, these models often fall short in incorporating relational information from critical epidemiological data sources such as human mobility, geographic connections, and contact tracing. This deficiency restricts their utility in broader epidemiological applications.
@@
Recently, the advances in Graph Neural Networks (GNNs) (SurveyonGNN_TNNLS20, 17, 18, 19, 20) have set the stage for overcoming the aforementioned challenges in epidemic modeling. Specifically, GNNs stand out for their ability to aggregate diverse information through a message-passing mechanism, making them particularly suited for capturing relational dynamics within graphs (liu2023uplifting, 21, 22). Thus, by representing interactions between entities as graphs, researchers can leverage GNNs to harness relational data effectively and facilitate epidemiology tasks (deng2020colagnn, 23, 24). For instance, GNNs are often utilized to model spatial interactions (gao2021stan, 25) and other complex interactions (caoMepoGNNMetapopulationEpidemic2022, 26), enhancing the analysis of graph data and yielding more precise predictions. In addition, GNNs bring a certain level of interpretability by quantifying the influence of individual nodes (or entities) on final prediction (feng2023contact, 27). Moreover, the flexible design of GNNs facilitates their integration with traditional mechanistic and probabilistic models to leverage expert knowledge and offer measures of uncertainty. As a result, GNNs have found extensive applications in various tasks within the field including infection prediction  (liu2023epidemiologyaware, 28) , outbreak source detection (ru2023inferring, 29), intervention modeling (song2020reinforced, 30), etc., facilitating advancements in epidemiology research.
@@
[ArXiv’21] (baldo2021deep, 16)
[JBD’21](shorten2021deep, 13)
[RBME’22](clement2022survey, 31)
[NC’22] (kamalov2022deep, 32)
[RBE’22] (dinesh10significance, 33)
[IJERPH’22] (saleem2022machine, 15)
[ArXiv’22] (Data-centric_arxiv22, 34)
@@
Connection to Existing Surveys. In contrast to previous surveys that explore the intersection of epidemiology and deep learning models, our paper offers a detailed overview specifically of GNNs in epidemic modeling. While preceding surveys predominantly concentrate on predicting COVID-19 outcomes and often overlook the inclusion of GNN-based methodologies (shorten2021deep, 13, 31, 33, 15), a few studies have indeed integrated GNNs (kamalov2022deep, 32, 16). However, the scope of such works remains relatively narrow, predominantly confined to virus transmission tasks. Certain studies concentrate exclusively on a singular virus (saleem2022machine, 15) or are dedicated to a specific task like forecasting (Data-centric_arxiv22, 34). Distinctively, our research is tailored towards GNN-based approaches, covering a broader spectrum of epidemic modeling tasks. Furthermore, our survey presents the latest review of GNN applications in epidemic modeling, offering deeper insights when compared to existing literature. The comparison can be found in Table 1. Figure 1 shows the structure of this survey.
@@ 2.2. Graph Neural Networks
Over recent years, GNNs have garnered increasing interest and have been deployed across diverse fields, including bioinformatics, material science, chemistry, and neuroscience (reiser2022graph, 35, 36, 37, 38). Among them, Graph Convolutional Networks (GCN) (GCN_ICLR17, 18) and Graph Attention Networks (GAT) (GAT_arxiv17, 19, 20), have advanced the frontier of research on graph-structured data with their sophisticated and effective designs (SurveyonGNN_TNNLS20, 17, 39, 40). Typically, GNNs aim to learn graph representations, including node embeddings $\textbf{h}_{i}\in\mathbb{R}^{d}$, by utilizing both the structural and feature information of a graph $G$. The process within a GNN involves two key operations: message passing and aggregation of neighborhood information. This involves each node in the graph repeatedly collecting and integrating information from its neighbors as well as its own attributes to enhance its representation. The operation of a GNN over $L$ layers can be described by the following expression:
@@ 2.3. Mechanistic Models
Empirical models (aleta2020modelling, 41, 42) in epidemic forecasting utilize historical data to discern patterns and forecast the future spread of diseases. In contrast, mechanistic models (jiang2021countrywide, 43, 44) provide a detailed framework that explores the biological and social complexities underlying the transmission of infectious diseases, thus exceeding the reliance on historical data inherent to empirical models. Among mechanistic approaches, classic compartmental models (1927contribution, 45) (e.g., SIR) are particularly notable. These models adeptly simplify the intricate dynamics of disease transmission into digestible components. This simplification facilitates a clearer understanding of infection spread, serving as a valuable tool for both researchers and policymakers.
SIR Compartmental Model. In the domain of epidemiology (danon2011networks, 46, 47), it is widely hypothesized that the rate at which networks evolve is significantly slower compared to the propagation speed of diseases. This fundamental assumption underpins the adoption of a SIR model (1927contribution, 45, 48), which is instrumental in accurately capturing the dynamics of epidemic spread. The SIR model categorizes the population into three distinct groups based on their disease status: susceptible (S) to infection, currently infectious (I), and recovered (R), with the latter group being immune to both contraction and transmission of the disease. The SIR model, formulated using ordinary differential equations (ODEs) (grassly2008mathematical, 49), are as follows:
@@
These equations distribute the total population $N$ across the aforementioned categories, with the transitions between states regulated by two pivotal parameters: the transmission rate $\beta$ ($S\to I$) and the recovery rate $\gamma$ ($I\to R$). The model posits a quadratic relationship for disease transmission via interactions between susceptible and infectious individuals ($\beta S(t)I(t)$), alongside a linear recovery mechanism ($\gamma I(t)$). By fine-tuning the parameters of the SIR model, it is possible to compute the basic reproduction number $R_{0}=\beta/\gamma$, serving as a metric for the disease transmission potentials (hanDevilLandscapesInferring2023, 50, 51).
SIR Variants. While the SIR model provides a powerful framework for analyzing disease dynamics, its simplicity can sometimes neglect critical factors such as incubation periods, non-permanent immunity, and heterogeneous population interactions. This limitation has spurred the development of SIR variants, which offer a more comprehensive and nuanced understanding of the spread and control of infectious diseases. We briefly outline some of the most commonly used variants: i) SEIR (dixon2018seds, 52): The SEIR model extends the basic SIR framework by incorporating an ’Exposed’ (E) compartment. This compartment represents individuals who have been exposed to an infectious disease but are not yet infectious themselves (1927contribution, 45, 53). The model details the transition through the stages according to the sequence: $S\rightarrow E\rightarrow I\rightarrow R$. ii) SIRD: Enhancing the traditional SIR model, the SIRD variant includes a ’Dead’ (D) compartment, thus adapting the progression to: $S\rightarrow I\rightarrow R\rightarrow D$. This modification accounts for individuals who succumb to the disease, providing a more accurate depiction of its mortality impact (tomyEstimatingStateEpidemics2022, 54, 55, 56).
@@
Detection SD-STGCN  (sha2021source, 57) GAT + GRU + SEIR Spatial-Temporal Graph; Static Graph Structure
Surveillance WDCIP  (wang2023wdcip, 58) GAE Spatial Graph; Static Graph Structure
Projection MMCA-GNNA (jhun2021effective, 59) GNN + SIR + RL Spatial-Temporal Graph; Static Graph Structure
DURLECA  (song2020reinforced, 30) GNN + RL
IDRLECA  (fengContactTracingEpidemic2023, 60) GNN + RL Spatial-Temporal Graph; Dynamic Graph Structure
Prediction DGDI  (liu2023human, 61) GCN + Self-Attention Spatial Graph; Static Graph Structure
DVGSN  (zhang2023predicting, 62) GNN Temporal-Only Graph; Static Graph Structure
STAN  (gao2021stan, 25) GAT + GRU Spatial-Temporal Graph; Static Graph Structure
MSDNet  (tang2023enhancing, 63) GAT + GRU + SIS
SMPNN  (lin2023grapha, 64) MPNN + Autoregression
ATMGNN  (nguyen2023predicting, 65) MPNN/MGNN + LSTM/Transformer
DASTGN  (puDynamicAdaptiveSpatio2023, 66) GNN + Attention + GRU
MSGNN  (qiu2023msgnna, 67) GCN + N-Beats
STEP  (yuSpatiotemporalGraphLearning2023, 68) GCN + Attention + GRU
GSRNN  (li2019study, 69) GNN + RNN
Mepo GNN (cao2023metapopulation, 70, 26) (TCN + GCN) + Modified SIR Spatial-Temporal Graph; Dynamic Graph Structure
Epi-Cola-GNN  (liu2023epidemiologyaware, 28) Cola-GNN + Modified SIR
CausalGNN  (wang2022causalgnn, 24) GNN + SIRD
ATGCN  (wang2022adaptively, 71) GNN + LSTM
HierST  (zheng2021hierst, 72) GNN + LSTM
RESEAT  (moon2023reseat, 73) GNN + Self-Attention
SAIFlu-Net  (jungSelfAttentionBasedDeepLearning2022, 74) GNN + LSTM
Epi-GNN  (xieEpiGNNExploringSpatial2022, 75) GCN + Attention + RNN
Cola-GNN  (deng2020colagnn, 23) GCN + Attention + RNN
@@ 3.1.1. Detection
For example, finding patient-zero  (ruInferringPatientZero2023, 76, 57, 51), as a source detection task, is important for identifying the source of disease outbreaks and aims to find a set of nodes $V$ on graph $G_{0}$. In this setting, the problem can also be seen as identifying the state of each node at the initial time point, which is $S_{0}^{V}$.

### v3

@@ Abstract.
Since the onset of the COVID-19 pandemic, there has been a growing interest in studying epidemiological models. Traditional mechanistic models mathematically describe the transmission mechanisms of infectious diseases. However, they often suffer from limitations of oversimplified or fixed assumptions, which could cause sub-optimal predictive power and inefficiency in capturing complex relation information. Consequently, Graph Neural Networks (GNNs) have emerged as a progressively popular tool in epidemic research. In this paper, we endeavor to furnish a comprehensive review of GNNs in epidemic tasks and highlight potential future directions. To accomplish this objective, we introduce hierarchical taxonomies for both epidemic tasks and methodologies, offering a trajectory of development within this domain. For epidemic tasks, we establish a taxonomy akin to those typically employed within the epidemic domain. For methodology, we categorize existing work into Neural Models and Hybrid Models. Following this, we perform an exhaustive and systematic examination of the methodologies, encompassing both the tasks and their technical details. Furthermore, we discuss the limitations of existing methods from diverse perspectives and systematically propose future research directions. This survey aims to bridge literature gaps and promote the progression of this promising field, with a list of relevant papers at https://github.com/Emory-Melody/awesome-epidemic-modeling-papers. We hope that it will facilitate synergies between the communities of GNNs and epidemiology, and contribute to their collective progress.
@@ 1. Introduction
Epidemiology has long been a critical field, with its origins tracing back to ancient societies that observed patterns of disease spread (bramanti2012ancient; bruce-chwatt1977plagues). Although the conceptualization of epidemiology has evolved over time, the terms health and control have been predominantly associated with it since 1978 (frerot2018epidemiology). Currently, the World Health Organization (WHO) describes epidemiology as the investigation into the distribution and determinants of health-related states or events, encompassing a broad spectrum of issues including disease transmission, vaccination efforts (fine2015another; terris1993society), cancer and diabetes treatment, etc. This definition underscores the field’s emphasis on controlling health-related issues and making informed decisions. A pertinent illustration of this is the COVID-19 pandemic, which rapidly infected millions worldwide, placing immense strain on the production and distribution of medical resources (cm2020does; bergrath2022impact). Decision-making processes like allocation of resources, are crucial in mitigating the impact of diseases and saving lives, highlighting the importance of advancements in epidemiology (emanuel2020fair).
@@
To address a range of health-related challenges, there is an indispensable need for epidemic modeling, and researchers have devised various mechanistic models (funk2018realtime; kondratyev2013forecasting). These models, grounded in mathematical formulations, simulate the dissemination of infectious diseases by incorporating biological and behavioral underpinnings. By considering factors such as population, they yield insights into patterns of disease transmission and the efficacy of intervention strategies, thereby playing a pivotal role in shaping public health policies (louz2010emergence; mikolajczyk2009influenza). However, these knowledge-driven methods often depend on oversimplified or fixed assumptions that can lead to biases in modeling. Consequently, this compromises both the accuracy of predictions and their ability to generalize across different contexts.
@@
To overcome the limitations of mechanistic models, there is an emerging trend to adopt data-driven approaches in epidemic forecasting tasks, with a particular emphasis on machine learning and deep learning models (shorten2021deep). Specifically, Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) have demonstrated great success in various epidemiological predictive tasks, including forecasting daily new case counts, estimating virus reproduction and doubling times, and determining disease-related factors (wu2018deep; saleem2022machine; baldo2021deep). Despite their effectiveness in these tasks, these models often fall short in incorporating relational information from critical epidemiological data sources such as human mobility, geographic connections, and contact tracing. This deficiency restricts their utility in broader epidemiological applications.
@@
Recently, the advances in Graph Neural Networks (GNNs) (SurveyonGNN_TNNLS20; GCN_ICLR17; GAT_arxiv17; GATv2_ICLR22) have set the stage for overcoming the aforementioned challenges in epidemic modeling. Specifically, GNNs stand out for their ability to aggregate diverse information through a message-passing mechanism, making them particularly suited for capturing relational dynamics within graphs (liu2023uplifting; maskey2022generalization). Thus, by representing interactions between entities as graphs, researchers can leverage GNNs to harness relational data effectively and facilitate epidemiology tasks (deng2020colagnn; wang2022causalgnn). For instance, GNNs are often utilized to model spatial interactions (gao2021stan) and other complex interactions (caoMepoGNNMetapopulationEpidemic2022), enhancing the analysis of graph data and yielding more precise predictions. In addition, GNNs bring a certain level of interpretability by quantifying the influence of individual nodes (or entities) on final prediction (feng2023contact). Moreover, the flexible design of GNNs facilitates their integration with traditional mechanistic and probabilistic models to leverage expert knowledge and offer measures of uncertainty. As a result, GNNs have found extensive applications in various tasks within the field including infection prediction  (liu2023epidemiologyaware) , outbreak source detection (ru2023inferring), intervention modeling (song2020reinforced), etc., facilitating advancements in epidemiology research.
@@
[ArXiv’21] (baldo2021deep)
[JBD’21](shorten2021deep)
[RBME’22](clement2022survey)
[NC’22] (kamalov2022deep)
[RBE’22] (dinesh10significance)
[IJERPH’22] (saleem2022machine)
[ArXiv’22] (Data-centric_arxiv22)
@@
Connections to Existing Surveys. In contrast to previous surveys that explore the intersection of epidemiology and deep learning models, our paper offers a detailed overview specifically of GNNs in epidemic modeling. While preceding surveys predominantly concentrate on predicting COVID-19 outcomes and often overlook the inclusion of GNN-based methodologies (shorten2021deep; clement2022survey; dinesh10significance; saleem2022machine), a few studies have indeed integrated GNNs (kamalov2022deep; baldo2021deep). However, the scope of such works remains relatively narrow, predominantly confined to virus transmission tasks. Certain studies concentrate exclusively on a singular virus (saleem2022machine) or are dedicated to a specific task like forecasting (Data-centric_arxiv22). Distinctively, our research is tailored towards GNN-based approaches, covering a broader spectrum of epidemic modeling tasks. Furthermore, our survey presents the latest review of GNN applications in epidemic modeling, offering deeper insights when compared to existing literature. The comparison can be found in Table 1. Figure 1 shows the structure of this survey.
@@ 2.2. Graph Neural Networks
Over recent years, GNNs have garnered increasing interest and have been deployed across diverse fields, including bioinformatics, material science, chemistry, and neuroscience (reiser2022graph; wen2022graph; wieder2020compact; bessadok2022graph). Among them, Graph Convolutional Networks (GCN) (GCN_ICLR17) and Graph Attention Networks (GAT) (GAT_arxiv17; GATv2_ICLR22), have advanced the frontier of research on graph-structured data with their sophisticated and effective designs (SurveyonGNN_TNNLS20; TrustworthyGNN_survey_arxiv22; deep_graph_clustering_survey_arxiv). Typically, GNNs aim to learn graph representations, including node embeddings $\textbf{h}_{i}\in\mathbb{R}^{d}$, by utilizing both the structural and feature information of a graph $G$. The process within a GNN involves two key operations: message passing and aggregation of neighborhood information. This involves each node in the graph repeatedly collecting and integrating information from its neighbors as well as its own attributes to enhance its representation. The operation of a GNN over $L$ layers can be described by the following expression:
@@ 2.3. Mechanistic Models
Empirical models (aleta2020modelling; chang2020modelling) in epidemic forecasting utilize historical data to discern patterns and forecast the future spread of diseases. In contrast, mechanistic models (jiang2021countrywide; yang2023epimob) provide a detailed framework that explores the biological and social complexities underlying the transmission of infectious diseases, thus exceeding the reliance on historical data inherent to empirical models. Among mechanistic approaches, classic compartmental models (1927contribution) (e.g., SIR) are particularly notable. These models adeptly simplify the intricate dynamics of disease transmission into digestible components. This simplification facilitates a clearer understanding of infection spread, serving as a valuable tool for both researchers and policymakers.
SIR Compartmental Model. In the domain of epidemiology (danon2011networks; caals2017ethics), it is widely hypothesized that the rate at which networks evolve is significantly slower compared to the propagation speed of diseases. This fundamental assumption underpins the adoption of a SIR model (1927contribution; dehning2020inferring), which is instrumental in accurately capturing the dynamics of epidemic spread. The SIR model categorizes the population into three distinct groups based on their disease status: susceptible (S) to infection, currently infectious (I), and recovered (R), with the latter group being immune to both contraction and transmission of the disease. The SIR model, formulated using ordinary differential equations (ODEs) (grassly2008mathematical), are as follows:
@@
These equations distribute the total population $N$ across the aforementioned categories, with the transitions between states regulated by two pivotal parameters: the transmission rate $\beta$ ($S\to I$) and the recovery rate $\gamma$ ($I\to R$). The model posits a quadratic relationship for disease transmission via interactions between susceptible and infectious individuals ($\beta S(t)I(t)$), alongside a linear recovery mechanism ($\gamma I(t)$). By fine-tuning the parameters of the SIR model, it is possible to compute the basic reproduction number $R_{0}=\beta/\gamma$, serving as a metric for the disease transmission potentials (hanDevilLandscapesInferring2023; shah2020finding).
SIR Variants. While the SIR model provides a powerful framework for analyzing disease dynamics, its simplicity can sometimes neglect critical factors such as incubation periods, non-permanent immunity, and heterogeneous population interactions. This limitation has spurred the development of SIR variants, which offer a more comprehensive and nuanced understanding of the spread and control of infectious diseases. We briefly outline some of the most commonly used variants: i) SEIR (dixon2018seds): The SEIR model extends the basic SIR framework by incorporating an ’Exposed’ (E) compartment. This compartment represents individuals who have been exposed to an infectious disease but are not yet infectious themselves (1927contribution; vandendriessche2017reproduction). The model details the transition through the stages according to the sequence: $S\rightarrow E\rightarrow I\rightarrow R$. ii) SIRD: Enhancing the traditional SIR model, the SIRD variant includes a ’Dead’ (D) compartment, thus adapting the progression to: $S\rightarrow I\rightarrow R\rightarrow D$. This modification accounts for individuals who succumb to the disease, providing a more accurate depiction of its mortality impact (tomyEstimatingStateEpidemics2022; wangCausalGNNCausalBasedGraph2022; lolipiccolomini2020monitoring).
@@
Detection SD-STGCN  (sha2021source) GAT + GRU + SEIR Spatial-Temporal Graph; Static Graph Structure
Surveillance WDCIP  (wang2023wdcip) GAE Spatial Graph; Static Graph Structure
Projection MMCA-GNNA (jhun2021effective) GNN + SIR + RL Spatial-Temporal Graph; Static Graph Structure
DURLECA  (song2020reinforced) GNN + RL
IDRLECA  (fengContactTracingEpidemic2023) GNN + RL Spatial-Temporal Graph; Dynamic Graph Structure
Prediction DGDI  (liu2023human) GCN + Self-Attention Spatial Graph; Static Graph Structure
DVGSN  (zhang2023predicting) GNN Temporal-Only Graph; Static Graph Structure
STAN  (gao2021stan) GAT + GRU Spatial-Temporal Graph; Static Graph Structure
MSDNet  (tang2023enhancing) GAT + GRU + SIS
SMPNN  (lin2023grapha) MPNN + Autoregression
ATMGNN  (nguyen2023predicting) MPNN/MGNN + LSTM/Transformer
DASTGN  (puDynamicAdaptiveSpatio2023) GNN + Attention + GRU
MSGNN  (qiu2023msgnna) GCN + N-Beats
STEP  (yuSpatiotemporalGraphLearning2023) GCN + Attention + GRU
GSRNN  (li2019study) GNN + RNN
Mepo GNN (cao2023metapopulation; caoMepoGNNMetapopulationEpidemic2022) (TCN + GCN) + Modified SIR Spatial-Temporal Graph; Dynamic Graph Structure
Epi-Cola-GNN  (liu2023epidemiologyaware) Cola-GNN + Modified SIR
HiSTGNN  (ma2022hierarchical) Hierarchical GNN + Transformer
CausalGNN  (wang2022causalgnn) GNN + SIRD
ATGCN  (wang2022adaptively) GNN + LSTM
HierST  (zheng2021hierst) GNN + LSTM
RESEAT  (moon2023reseat) GNN + Self-Attention
SAIFlu-Net  (jungSelfAttentionBasedDeepLearning2022) GNN + LSTM
Epi-GNN  (xieEpiGNNExploringSpatial2022) GCN + Attention + RNN
Cola-GNN  (deng2020colagnn) GCN + Attention + RNN
@@ 3.1.1. Detection
For example, finding patient-zero  (ruInferringPatientZero2023; sha2021source; shah2020finding), as a source detection task, is important for identifying the source of disease outbreaks and aims to find a set of nodes $V$ on graph $G_{0}$. In this setting, the problem can also be seen as identifying the state of each node at the initial time point, which is $S_{0}^{V}$.